OpenStudy (anonymous):

What is the slope of the line 5x + 2y = 10?

hero (hero):

To find the slope of any line, manipulate the equation and re-write the line so that it is in the form y = mx + b where m = slope. In this particular case, To find the slope of 5x + 2y = 10, manipulate and rewrite it as follows: 5x + 2y = 10 2y = 10 - 5x y = (10 - 5x)/2 y = 5 - 2.5x y = -2.5x + 5 Thus now, the line 5x + 2y = 10 re-written as y = -2.5x + 5 is in the form y = mx + b and m, the slope = -2.5